Titanium is a transition metal used in many alloys because it is extremely strong and lightweight. Titanium tetrachloride is extracted from titanium oxide using chlorine and coke. What mass of C is needed to react with 1.25 mol of Tio2?


1
Expert's answer
2022-05-04T09:27:03-0400

The titanium extraction from its ore is done by chloride and coke by the following equation:


"TiO_2+2Cl_2+2C= TiCl_4 + 2CO"

By the reaction, the number of C moles required to extract titanium is two times more as compared to "TiO_2", thus the mol number of C is:


"n(C)=2\times n(TiO_2) = 2\times 1.25 mol = 2.5 mol"

The mass of carbon is:


"m(C) = n(C) \times M(C) = 2.5 mol \times 12 g/mol = 30 g"